solve for the positive value of x: 6x^2-54=0
Answer by jim_thompson5910(35256)

6x^2-54=0

6x^2-54 + 54=0+54

6x^2=54

(6x^2)/6=54/6

x^2 = 9

sqrt(x^2) = sqrt(9)

x = 3


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim